Give an example of a nowhere dense set in a metric space .substantiate your claim


1
Expert's answer
2021-07-19T08:10:27-0400

Solution:

Let X be a metric space. A subset A ⊆ X is called nowhere dense in X if the interior of the closure of A is empty, i.e. (A) = ∅. Otherwise put, A is nowhere dense iff it is contained in a closed set with empty interior.

For example, "\mathbb{Z}" is nowhere dense in "\mathbb{R}" because it is its own closure, and it does not contain any open intervals (i.e. there is no (a, b) s.t. "(a, b) \subset \overline{\mathbb{Z}}=\mathbb{Z}" . An example of a set which is not dense, but which fails to be nowhere dense would be "\{x \in \mathbb{Q} \mid 0<x<1\}" . Its closure is [0,1], which contains the open interval (0,1). Using the alternate definition, you can note that the set is dense in "(0,1) \subset \mathbb{R}" .